OpenStudy (anonymous):

5 + x = x + 5 what property? Identity Property for Multiplication Inverse Property for Addition Commutative for Addition Distributive Property

OpenStudy (accessdenied):

identity prop. of multiplication is \(a*1 = a\) inverse prop. of addition is \(a + (-a) = 0\) commutative prop. of addition is \(a + b = b + a\) distributive prop. is \(a(b + c) = ab + ac\) Which does it most resemble?

OpenStudy (anonymous):

commutative prop?

OpenStudy (accessdenied):

Correct. :)